OpenStudy (anonymous):

What is 8.3 repeating as a fraction ?

OpenStudy (aum):

I answered this when you posted before. If you want a different method here it is: 8.33333..... = 8 + .33333.... = 8 + 1/3 = 25/3 (because we know 1/3 = .3333..... (repeating))